# Probing the accretion geometry of black hole X-ray binaries: A multi-mission spectro-polarimetric and timing study

### v1

@@ Abstract
We present a comprehensive spectro-polarimetric and timing analysis of eleven black hole X-ray binaries, namely Cyg X11-1- 1, 4U $1630-47$, Cyg X33-3- 3, LMC X11-1- 1, 4U $1957+115$, LMC X33-3- 3, Swift J1727.816131727.816131727.8-16131727.8 - 1613, GX $339-4$, Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147, IGR J17091362417091362417091-362417091 - 3624, and MAXI J174429417442941744-2941744 - 294, using quasi-simultaneous observations from IXPE, NICER, NuSTAR, and AstroSat. Timing analyses reveal characteristic type-B and type-C Quasi-periodic Oscillations (QPOs) across different spectral states, often associated with episodic radio ejections. Broadband ($0.7-60$ keV) spectral modelling, employing disc, Comptonization, and reflection components, reveals degeneracies in constraining disc-corona geometries. Polarimetric measurements in $2-8$ keV band detect significant polarization degrees (PDs) ranging from $3-20.6\%$ ($1.2-21.4\%$) in harder (softer) states, with moderate to strong energy dependence, except for LMC X11-1- 1, Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147, and MAXI J174429417442941744-2941744 - 294, where no significant polarization is detected. We report the polarization detections of Cyg X33-3- 3 (PD $\sim 21.4\%$, SIMS), LMC X33-3- 3 (PD $\sim 2.4\%$, HSS) and IGR J17091362417091362417091-362417091 - 3624 (PD $\sim 9\%$, LHS) using the recent IXPE observations. A positive correlation is found between PD and the Comptonized photon fraction ($cov_{\rm frac}$), while an anti-correlation is observed with the disc-to-Comptonized flux ratio ($F_{\rm ratio}$) across spectral states. The combined timing, spectral, and polarimetric results, together with constraints from radio jet observations, suggest a radially extended corona within a truncated disc for Cyg X11-1- 1, Swift J1727.816131727.816131727.8-16131727.8 - 1613, IGR J17091362417091362417091-362417091 - 3624, and GX $339-4$, whereas the disc-corona geometry remains unconstrained for 4U $1957+115$, LMC X33-3- 3, and 4U $1630-47$. We discuss the implications of these findings for understanding accretion geometries and highlight prospects for future X-ray polarimetric studies.
@@ 1 Introduction
Black hole X-ray binaries (BH-XRBs) are believed to be the ideal candidates for understanding the physical processes that govern the radiation mechanism around the compact objects. BH-XRBs often exhibit distinct spectral states over timescales ranging from days to months, strongly associated with the underlying accretion dynamics of the system (Morgan et al., 1997; Zhang et al., 1997; Paul et al., 1998; Sobczak et al., 1999; Done & Zycki, 1999; Chakrabarti & Manickam, 2000; Homan et al., 2001; Corbel et al., 2001; Titarchuk & Shrader, 2002; Pottschmidt et al., 2003; Rodriguez et al., 2003; Corbel et al., 2004; Belloni et al., 2005; Kalemci et al., 2006; Remillard & McClintock, 2006; Yu & Yan, 2009; Shaposhnikov et al., 2010; Tarana et al., 2011; Nandi et al., 2012; Sriram et al., 2013; Seifina et al., 2014; Iyer et al., 2015; Yan & Yu, 2017; Sreehari et al., 2019b, 2020; Baby et al., 2020; Majumder et al., 2022; Athulya et al., 2022; Prabhakar et al., 2023; Nandi et al., 2024; Li et al., 2025; Majumder et al., 2025b, and references therein). Empirically, the high-soft state (HSS) is characterized by the multi-temperature black-body emission that is likely to emerge from an optically thick and geometrically thin accretion disc Shakura & Sunyaev (1973). In contrast, the canonical low-hard state (LHS), represented by a power-law profile with high energy cut-off, is found to be dominated by the hard emission produced from the Compton upscattering of disc photons into the ‘hot’ electron cloud (equivalently X-ray corona). Moreover, depending on the relative contributions from the disc and coronal components, hard/soft intermediate states (HIMS/SIMS) are observed.
@@
Notably, the temporal properties of BH-XRBs show rapid X-ray variability over different timescales. This variability is usually observed in the power density spectrum (PDS) and is closely correlated to the spectral states. In particular, transient phenomena like strong and stable Low-Frequency Quasi-periodic Oscillations (LFQPOs) on a wide range of frequencies distinguish the spectral states and act as the precursor of the state transitions in BH-XRBs (Remillard & McClintock, 2006; Done et al., 2007; Nandi et al., 2012; Iyer et al., 2015). For example, the LHS and HIMS are characterized by the appearance of strong, coherent, variable peaked type-C LFQPO of frequency $\sim 0.1-15$ Hz superposed on a flat-top noise (FTN) component in the PDS (Remillard & McClintock, 2006; Nandi et al., 2012, and references therein). The origin of these type-C QPOs is often explained through various mechanisms, including oscillations of radiative shock waves within the accretion disc (Molteni et al., 1996; Chakrabarti et al., 2008; Das et al., 2014), relativistic Lense-Thirring precession of the inner hot flow or the truncated disc (Stella & Vietri, 1998; Ingram et al., 2009), precession of small-scale jets (Ma et al., 2021) and outward drift of the truncated inner disc radius enveloped by the corona (Karpouzas et al., 2020; Bellavita et al., 2022, 2025). The wide array of interpretations introduces a degeneracy among the different disc-corona-jet configurations used in explaining the QPO phenomena (see Ingram & Motta 2019 for a review).
@@
Furthermore, X-ray polarimetric study is also considered as a powerful diagnostic tool to infer the accretion geometry of the BH-XRBs. The recent launch of IXPE Weisskopf et al. (2022), a polarimetric mission sensitive to low-energy ($2-8$ keV) X-rays, enables the opportunity to investigate in-depth polarimetric properties of BH-XRBs (see Dovčiak et al. 2024 for a summary). So far, IXPE has observed eleven BH-XRBs, with significant polarized emission detected in nine sources, namely Cyg X11-1- 1 Krawczynski et al. (2022); Jana & Chang (2024); Steiner et al. (2024), 4U $1630-47$ Kushwaha et al. (2023b); Rawat et al. (2023a, b); Rodriguez Cavero et al. (2023); Ratheesh et al. (2024), Cyg X33-3- 3 Veledina et al. (2024a, b), LMC X33-3- 3 Majumder et al. (2024a); Svoboda et al. (2024a), 4U $1957+115$ Kushwaha et al. (2023a); Marra et al. (2024), Swift J1727.816131727.816131727.8-16131727.8 - 1613 Veledina et al. (2023); Ingram et al. (2024); Svoboda et al. (2024b); Podgorný et al. (2024), Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147 Mondal et al. (2024), GX $339-4$ Mastroserio et al. (2025) and IGR J17091362417091362417091-362417091 - 3624 Ewing et al. (2025). Notably, no significant polarization is observed in LMC X11-1- 1 Podgorný et al. (2023).
@@
Despite the significant advancements in X-ray polarimetry, interpreting the observed polarization degree (PD) and polarization angle (PA) within the framework of theoretical models remains a formidable challenge. The classical work by Chandrasekhar (Chandrasekhar, 1960) on semi-infinite electron scattering predicts a low PD of $\sim 2\%$ from the accretion disc of highly inclined systems. However, 4U 1630--47 shows a remarkably higher PD of around $8.3\%$ in its disc-dominated thermal state, deviating substantially from these theoretical expectations (Kushwaha et al., 2023b; Ratheesh et al., 2024). Furthermore, PDs of about $4\%$ have been observed in the LHS of Cyg X--1 and Swift J1727.8--1613 (Krawczynski et al., 2022; Ingram et al., 2024), whereas existing models predict only $\sim 1\%$ PD from a wedge-shaped corona in such low-inclined systems (Krawczynski et al., 2022). To reconcile this discrepancy, Poutanen et al. (2023) proposed that an outflowing corona with mildly relativistic motion could produce PD levels consistent with observations. It is worth mentioning that the recent detection of $\sim 9\%$ PD in the LHS of IGR J17091--3624 (Ewing et al., 2025) further indicates the limitations of current theoretical frameworks and highlights the need for more comprehensive models of X-ray polarization.
@@
Moreover, in certain cases, the spectro-temporal findings of BH-XRBs appear to be contradictory with interpretations derived from X-ray polarimetry. For instance, in Swift J1727.8--1613, the evolution of type-C QPOs suggests the presence of a jet-like corona aligned perpendicular to the disc plane (Liao et al., 2024). In contrast, simultaneous polarimetric observations of the same source indicate a radially extended corona resided at the equatorial plane of the disc (Veledina et al., 2023; Ingram et al., 2024). Therefore, it is evident that the polarimetric findings offer challenges to the existing X-ray spectro-temporal models, indicating that a deeper understanding of the accretion dynamics and disc-corona-jet geometry in black hole X-ray binaries is yet to be unveiled.
@@
Being motivated by this, we undertake in-depth timing and spectro-polarimetric analyses of eleven BH-XRBs using quasi-simultaneous archival observations from IXPE, NICER, NuSTAR and AstroSat, covering a broad energy range of $0.5-100$ keV. While doing so, we study the evolution of X-ray (soft/hard) and radio flux densities for all sources using multi-mission long-term monitoring data from MAXI/GSC, Swift/BAT, RATAN and VLITE. Further, we conduct a comprehensive spectro-temporal study of these sources using NICER, NuSTAR and AstroSat data and confirm the presence of distinct spectral states of the BH-XRBs during the observational campaigns. Using alternative model prescriptions relying on different assumptions on the disc-corona configurations, we attempt to explain the spectral features of the sources and investigate the possibility of having inherent degeneracy among the models. Next, we deduce the polarization properties of the sources using IXPE data in the $2-8$ keV energy range, followed by a detailed spectro-polarimetric correlation study. Subsequently, noticeable positive and negative correlations between polarization degree and spectral parameters serve as powerful diagnostics, offering deeper insights into the complex accretion dynamics and geometry of the BH-XRBs under consideration. Finally, we outline the future prospects of X-ray polarimetry in light of the present findings.
@@
Figure 1: Long-term X-ray monitoring of the sources with MAXI/GSC ($2-20$ keV), Swift/BAT ($15-50$ keV) and the corresponding hardness ratio obtained with MAXI/GSC are presented in panel (a), (b) and (c), respectively. Each data point for a given source represents averaged out respective quantity over the exposure of corresponding IXPE observation. In panel (d), the variation of radio flux is shown from the quasi-simultaneous radio observations available from various facilities. The $1\sigma$ errorbars of the respective quantities are small and remain within the markers for most of the cases. For a given source, the data points plotted with MJD, correspond to the respective epochs as mentioned in Table 1. See the text for details.
@@ 2 Source Selection and observations
In this work, we analyze all the IXPE (Weisskopf et al., 2022) observations of BH-XRBs conducted so far along with simultaneous/quasi-simultaneous NICER, NuSTAR and AstroSat observations depending on data availability. As of now, IXPE observed eleven BH-XRBs, namely Cyg X11-1- 1, 4U 1630--47, Cyg X33-3- 3, LMC X11-1- 1, 4U 1957$+115$, LMC X33-3- 3, Swift J1727.816131613-1613- 1613, Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147, GX $339-4$, IGR J17091362417091362417091-362417091 - 3624 and MAXI J174429417442941744-2941744 - 294, during its first three and a half years of operational period. The IXPE, NICER and NuSTAR observations used in this work are publicly available in the HEASARC [^1] https://heasarc.gsfc.nasa.gov/db-perl/W3Browse/w3browse.pl database and the AstroSat data is archived at the ISSDC [^2] https://webapps.issdc.gov.in/astro_archive/archive/Home.jsp website. The quasi-simultaneous multi-mission observations of the sources are separated by at most three days from the IXPE epoch except LMC X11-1- 1 for which the NuSTAR observation lies four days after the IXPE observation. The sources exhibiting marginal spectro-temporal variability between observations from different missions offer an opportunity for a combined multi-mission study. All the observations analyzed in this study are tabulated in Table 1 along with the date and ObsID. Brief overview of each of the sources under consideration are presented below.
2.1 Cyg X--1
@@
Cyg X-1, the first galactic BH-XRB discovered in 1971 Webster & Murdin (1972), continues to be one of the most extensively studied celestial objects. Recent measurements reveal that it hosts a black hole with a mass of $21.2\pm 2.2$ $M_{\odot}$ Miller-Jones et al. (2021), located at a distance of $2.2\pm 0.2$ kpc Miller-Jones et al. (2021), with a supergiant O-type star as its binary companion. Meanwhile, numerous studies confirm the presence of a maximally rotating ‘hole’ having spin greater than $0.99$ at the center core of the binary system (Tomsick et al., 2013; Zhao et al., 2021; Kushwaha et al., 2021). Cyg X--1 has remained persistently bright, mostly in the LHS, though it has transitioned to the HSS through short-lived intermediate states over the past few decades (Kushwaha et al., 2021).
2.2 4U 1630--47
@@
4U 1630--47 is a recurrent X-ray transient discovered by Uhuru Jones et al. (1976). Since discovery, the source has exhibited more than $20$ quasi-period outbursts Baby et al. (2020); Chatterjee et al. (2022) typically in every $600-700$ days Parmar et al. (1995). Since the optical counterpart of the source is still unknown, the dynamical measurement of its mass and distance remains elusive. Several efforts relying on indirect measurements constrained its mass as $10\pm 0.1$ $M_{\odot}$ Seifina et al. (2014) and distance as $10-11$ kpc Seifina et al. (2014); Kalemci et al. (2018). The presence of dips in the light curves of the $1996$ outburst indicates a high inclination ($i\sim 60-70^{\circ}$) of the system Tomsick et al. (1998); Seifina et al. (2014). Spectral modeling Pahari et al. (2018) and spectro-polarimetric fitting in the high soft state confirm the source’s spin as $\sim 0.92$ Kushwaha et al. (2023b). 4U 16304747-47- 47 is observed in the HSS in most of its outbursts till date, possibly due to fast transition from LHS to HSS Capitanio et al. (2015); Baby et al. (2020); Peng et al. (2024c); Fan et al. (2024); Parra et al. (2025).
2.3 Cyg X--3
@@
Cyg X33-3- 3 was discovered over five decades ago Giacconi et al. (1967) and is a high-mass X-ray binary system hosting a compact object accreting from a Wolf-Rayet donor star van Kerkwijk et al. (1992). Despite extensive studies over the years, the nature of the compact object remains uncertain. Recent polarization measurements suggest that it could be a concealed Galactic ultra-luminous X-ray source (ULX) surpassing the Eddington limit due to anisotropic emissions Veledina et al. (2024a). It is worth mentioning that Cyg X33-3- 3 exhibits exceptional radio activity Martí et al. (2000); Miller-Jones et al. (2004) which is tightly correlated with the spectral states on several occasions including the detection of innermost jet confined within the X-ray funnel-like structure Yang et al. (2023). Recent trigonometric parallax measurements from VLBA observations have precisely constrained the source’s distance as ${9.67}_{-0.48}^{+0.53}$ kpc Reid & Miller-Jones (2023).
2.4 LMC X--1
@@
LMC X11-1- 1 is the first discovered extra-galactic BH-XRB located in the Large Magellanic Cloud (LMC) Mark et al. (1969) at a well-constrained distance of $\sim 48.1$ kpc Pietrzyński et al. (2013). Dynamical measurements have determined the black hole’s mass to be $10.9\pm 1.4$ $M_{\odot}$ with a moderate inclination of $36.4^{{}^{\circ}}\pm 1.9^{{}^{\circ}}$ Orosz et al. (2009). Unlike typical X-ray binaries, LMC X11-1- 1 consistently remains in the HSS, with thermal emission contributing around $80\%$ of the spectral flux Steiner et al. (2012); Bhuvana et al. (2021); Jana et al. (2021) and a luminosity of approximately $2\times 10^{38}$ erg $\rm s^{-1}$ Nowak et al. (2001); Bhuvana et al. (2021). The measurements using continuum fitting method in the thermally dominated state constraints LMC X11-1- 1 spin as $\sim 0.92$ Gou et al. (2009); Bhuvana et al. (2021), indicating the presence of a rapidly rotating black hole.
2.5 4U 1957$+$115
4U 1957$+$115 is a bright and persistent low-mass X-ray binary, first discovered by the Uhuru mission in $1973$ Giacconi et al. (1974). To date, it has consistently been observed in a disc-dominated spectral state, with flux levels varying between $20$ and $70$ mCrab Yaqoob et al. (1993); Maitra et al. (2014); Mudambi et al. (2022). Despite decades of observations with various X-ray missions, key system parameters such as distance, inclination, spin, and mass remain highly uncertain. For instance, a high source distance of $20-40$ kpc and a low inclination ($\sim 13^{\circ}$) have been proposed for a maximum black hole mass of $\sim 6$ $M_{\odot}$ Gomez et al. (2015). In contrast, Maitra et al. (2014) reported a high inclination ($\sim 78^{\circ}$) and a distance between $5-10$ kpc for a black hole mass of less than $10$ $M_{\odot}$. Nevertheless, several studies suggest the presence of a maximally spinning black hole in 4U 1957$+$115 Barillier et al. (2023).
2.6 LMC X--3
@@
LMC X33-3- 3 is a persistent extra-galactic BH-XRB in Large Magellanic Cloud (LMC) at a distance of $48.1$ kpc Orosz et al. (2009), discovered by UHURU in $1971$. A slowly rotating black hole of spin $0.25_{-0.29}^{+0.20}$ Steiner et al. (2010); Bhuvana et al. (2021) and mass $6.98\pm 0.56~{}{\rm M}_{\odot}$ Orosz et al. (2014); Bhuvana et al. (2021) is found to be present at the central core of the system. LMC X33-3- 3 is a relatively high inclined system ($i\sim 69.24^{\circ}\pm 0.72^{\circ}$) Orosz et al. (2014). Being a persistent source, it mostly remains in the HSS Bhuvana et al. (2021) except for a few occasions during which the LHS is also observed Smale & Boyd (2012). An exceptionally anomalous low state of peak luminosity $\sim 10^{35}$ erg $\rm s^{-1}$ is also observed for a few instances Torpin et al. (2017).
2.7 Swift J1727.8--1613

### v2

@@ Abstract
We present a comprehensive spectro-polarimetric and timing analysis of twelve black hole X-ray binaries, namely Cyg X-1, 4U 1630-47, Cyg X-3, LMC X-1, 4U 1957++115, LMC X-3, Swift J1727.8-1613, GX 339-4, Swift J151857.0-572147, IGR J17091-3624, MAXI J1744-294 and GRS 1915++105, using quasi-simultaneous observations from IXPE, NICER, NuSTAR, and AstroSat. Timing analyses reveal type-B and type-C Quasi-periodic Oscillations across different spectral states, often associated with episodic radio ejections. Broadband (0.7600.7-60 keV) spectral modeling, employing disc, Comptonization, and reflection components, reveals degeneracies in constraining disc-corona geometries. Polarimetric measurements in 282-8 keV band detect significant polarization degrees (PDs) ranging from $3-20.6\%$ ($1.2-21.4\%$) in harder (softer) states, with moderate to strong energy dependence, except for LMC X1-1, Swift J151857.0572147151857.0-572147, and MAXI J1744-294, where no significant polarization is detected. We report polarization detections of Cyg X3-3 (PD $\sim 21.4\%$, SIMS), LMC X3-3 (PD $\sim 2.4\%$, HSS) and IGR J17091362417091-3624 (PD $\sim 9\%$, LHS) using recent IXPE observations. A positive correlation is found between PD and Comptonized photon fraction ($cov_{\rm frac}$), while an anti-correlation is observed with disc-to-Comptonized flux ratio ($F_{\rm ratio}$) across spectral states. Combined timing, spectral, and polarimetric results, together with constraints from radio jet observations, suggest a radially extended corona within a truncated disc for Cyg X1-1, Swift J1727.816131727.8-1613, IGR J17091362417091-3624, and GX 3394339-4, whereas the disc-corona geometry remains poorly constrained for 4U $1957+115$, LMC X3-3, and 4U 1630471630-47. We discuss the implications of these findings for understanding accretion geometries and highlight prospects for future X-ray polarimetric studies.
@@ 1 Introduction
Black hole X-ray binaries (BH-XRBs) are believed to be the ideal candidates for understanding the physical processes that govern the radiation mechanism around the compact objects. BH-XRBs often exhibit distinct spectral states over timescales ranging from days to months, strongly associated with the underlying accretion dynamics of the system (Morgan et al., 1997; Done & Zycki, 1999; Chakrabarti & Manickam, 2000; Homan et al., 2001; Belloni et al., 2005; Remillard & McClintock, 2006; Nandi et al., 2012; Iyer et al., 2015; Sreehari et al., 2019b; Baby et al., 2020; Majumder et al., 2022; Athulya et al., 2022; Prabhakar et al., 2023; Nandi et al., 2024; Majumder et al., 2025b, and references therein). Empirically, the high-soft state (HSS) is characterized by the multi-temperature black-body emission that is likely to emerge from an optically thick and geometrically thin accretion disc Shakura & Sunyaev (1973). In contrast, the canonical low-hard state (LHS), represented by a power-law profile with high energy cut-off, is found to be dominated by the hard emission produced from the Compton upscattering of disc photons into the ‘hot’ electron cloud (equivalently X-ray corona). Moreover, depending on the relative contributions from the disc and coronal components, hard/soft intermediate states (HIMS/SIMS) are observed.
@@
Notably, the temporal properties of BH-XRBs show rapid X-ray variability over different timescales. This variability is usually observed in the power density spectrum (PDS) and is closely correlated to the spectral states. In particular, transient phenomena like strong and stable Low-Frequency Quasi-periodic Oscillations (LFQPOs) on a wide range of frequencies distinguish the spectral states and act as the precursor of the state transitions in BH-XRBs (Remillard & McClintock, 2006; Done et al., 2007; Nandi et al., 2012; Iyer et al., 2015). For example, the LHS and HIMS are characterized by the appearance of strong, coherent, variable peaked type-C LFQPO of frequency $\sim 0.1-15$ Hz superposed on a flat-top noise (FTN) component in the PDS (Remillard & McClintock, 2006; Nandi et al., 2012, and references therein). The origin of these type-C QPOs is often explained through various mechanisms, including oscillations of radiative shock waves within the accretion disc (Molteni et al., 1996; Chakrabarti et al., 2008; Das et al., 2014), relativistic Lense-Thirring precession of the inner hot flow or the truncated disc (Stella & Vietri, 1998; Ingram et al., 2009), precession of small-scale jets (Ma et al., 2021) and outward drift of the truncated inner disc radius enveloped by the corona (Karpouzas et al., 2020; Bellavita et al., 2022, 2025). The wide array of interpretations introduces a degeneracy among different disc-corona-jet configurations used in explaining QPO phenomena (see Ingram & Motta 2019 for a review).
@@
Furthermore, X-ray polarimetric study is also considered a powerful diagnostic tool to infer the accretion geometry of the BH-XRBs. The recent launch of IXPE Weisskopf et al. (2022), a polarimetric mission sensitive to low-energy (282-8 keV) X-rays, enables the opportunity to investigate in-depth polarimetric properties of BH-XRBs. So far, IXPE has observed twelve BH-XRBs, namely Cyg X-1, 4U 1630-47, Cyg X-3, LMC X-3, LMC X-1, 4U 1957++115, Swift J1727.8-1613, Swift J151857.0-572147, GX 339-4, IGR J17091-3624, MAXI J1744-294 and GRS 1915++105 with significant polarized emission detected in eight sources (see also Dovčiak et al., 2024).
@@
Despite the significant advancements in X-ray polarimetry, interpreting the observed polarization degree (PD) and polarization angle (PA) within the framework of theoretical models remains a formidable challenge. The classical work by Chandrasekhar (Chandrasekhar, 1960) on semi-infinite electron scattering predicts a low PD of $\sim 2\%$ from the accretion disc of highly inclined systems. However, 4U 1630-47 shows a remarkably higher PD of around $8.3\%$ in its disc-dominated thermal state, deviating substantially from these theoretical expectations (Kushwaha et al., 2023; Ratheesh et al., 2024). Similarly, soft state observations of Cyg X1-1 show polarization perpendicular to the disc plane, whereas theoretical models predict it to be parallel. This possibly indicates a rapidly rotating black hole in Cyg X1-1, responsible for the returning radiation effects seen in the spectro-polarimetric data (Steiner et al., 2024; Krawczynski et al., 2024). Furthermore, PDs of about $4\%$ have been observed in the LHS of Cyg X-1 and Swift J1727.8-1613 (Krawczynski et al., 2022; Ingram et al., 2024), whereas existing models predict only $\sim 1\%$ PD from a wedge-shaped corona in such low-inclined systems (Krawczynski et al., 2022). To reconcile this discrepancy, Poutanen et al. (2023) invoked the model of Beloborodov (1998) of an outflowing corona with mildly relativistic motion to explain the high PD levels. It is worth mentioning that the recent detection of $\sim 9\%$ PD in the LHS of IGR J17091-3624 (Ewing et al., 2025) further indicates the limitations of current theoretical frameworks and highlights the need for more comprehensive models of X-ray polarization.
@@
Moreover, in certain cases, the spectro-temporal findings of BH-XRBs appear to be contradictory with interpretations derived from X-ray polarimetry. For instance, in Swift J1727.8-1613, the evolution of type-C QPOs suggests the presence of a jet-like corona aligned perpendicular to the disc plane (Liao et al., 2024). In contrast, simultaneous polarimetric observations of the same source indicate a radially extended corona resided at the equatorial plane of the disc (Veledina et al., 2023; Ingram et al., 2024). Therefore, it is evident that the polarimetric findings offer challenges to the existing X-ray spectro-temporal models, indicating that a deeper understanding of the accretion dynamics and disc-corona-jet geometry in BH-XRBs is yet to be unveiled.
@@
Being motivated by this, we carry out in-depth timing and spectro-polarimetric analyses of eleven BH-XRBs using quasi-simultaneous archival observations from IXPE, NICER, NuSTAR and AstroSat, covering a broad energy range of 0.51000.5-100 keV. We conduct a comprehensive spectro-temporal study of these sources using NICER, NuSTAR and AstroSat data and confirm the presence of distinct spectral states of the BH-XRBs during the observational campaigns. Further, we deduce the polarization properties of the sources using IXPE data in the 282-8 keV energy range, followed by a detailed spectro-polarimetric correlation study. Notably, the timing and spectro-polarimetric results serve as powerful diagnostics offering deeper insights into the complex accretion dynamics and geometry of BH-XRBs under consideration.
@@
Figure 1: Long-term X-ray monitoring of the sources with MAXI/GSC (2202-20 keV), Swift/BAT (155015-50 keV) and the corresponding hardness ratio obtained with MAXI/GSC are presented in panel (a), (b) and (c), respectively. Each data point for a given source represents averaged out respective quantity over the exposure of corresponding IXPE observation. In panel (d), the variation of radio flux is shown from the quasi-simultaneous radio observations available from various facilities. The $1\sigma$ errorbars of the respective quantities are small and remain within the markers for most of the cases. For a given source, the data points plotted with MJD, correspond to the respective epochs as mentioned in Table 1.
@@ 2 Source Selection and observations
In this work, we analyze all the IXPE (Weisskopf et al., 2022) observations of BH-XRBs conducted so far along with simultaneous/quasi-simultaneous NICER, NuSTAR and AstroSat observations depending on data availability. As of now, IXPE observed twelve BH-XRBs during its first three and a half years of operational period. The IXPE, NICER and NuSTAR observations used in this work are publicly available in the HEASARC [^1] https://heasarc.gsfc.nasa.gov/db-perl/W3Browse/w3browse.pl database and the AstroSat data is archived at the ISSDC [^2] https://webapps.issdc.gov.in/astro_archive/archive/Home.jsp website. The quasi-simultaneous multi-mission observations of the sources are separated by at most three days from the IXPE epoch except LMC X1-1 for which the NuSTAR observation lies four days after the IXPE observation. The sources exhibiting marginal spectro-temporal variability between observations from different missions offer an opportunity for a combined multi-mission study. We mention that IXPE monitoring of GRS $1915+105$ was carried out on $10^{\rm th}$ May 2025 for about $\sim 142$ ks. However, the source remained below the detection level with IXPE, showing a MAXI/GSC flux of $\sim 6$ mCrab (see Fig. 1). Therefore, GRS $1915+105$ is excluded from the detailed analyses in the following sections. All the observations analyzed in this study are tabulated in Table 1. A brief overview of each of the sources under consideration are presented below.
@@
$\bullet$ Cyg X-1: This is the first galactic BH-XRB discovered in 1971 Webster & Murdin (1972), continues to be one of the most extensively studied celestial objects. Recent measurements reveal that it hosts a black hole with a mass of $21.2\pm 2.2$ $M_{\odot}$, located at a distance of $2.2\pm 0.2$ kpc, with a supergiant O-type star as its binary companion (Miller-Jones et al., 2021). Meanwhile, numerous studies confirm the presence of a maximally rotating black hole having spin greater than $0.99$ at the center core of the binary system (Tomsick et al., 2013; Gou et al., 2014; Zhao et al., 2021; Kushwaha et al., 2021). Cyg X-1 has remained persistently bright, mostly in the LHS, though it has transitioned to the HSS through short-lived intermediate states over the past few decades (Kushwaha et al., 2021). Interestingly, significant polarized emission is detected in both LHS and HSS of the source (Krawczynski et al., 2022; Jana & Chang, 2024; Steiner et al., 2024).
@@
$\bullet$ 4U 1630-47: This is a recurrent X-ray transient discovered by Uhuru Jones et al. (1976). Since discovery, the source has exhibited more than $20$ quasi-period outbursts Baby et al. (2020); Chatterjee et al. (2022) typically in every 600700600-700 days Parmar et al. (1995). Several efforts relying on indirect measurements constrained its mass as $10\pm 0.1$ $M_{\odot}$ Seifina et al. (2014), distance as 101110-11 kpc Seifina et al. (2014); Kalemci et al. (2018) and high inclination $i\sim 60-70^{\circ}$ Tomsick et al. (1998); Seifina et al. (2014). Spectral modeling Pahari et al. (2018) and spectro-polarimetric fitting in the high soft state confirm its spin as $\sim 0.92$ Kushwaha et al. (2023). Notably, 4U 1630-47 is the second source observed by IXPE for which a remarkably high polarization degree ($\sim 6-8\%$) has been detected in both thermal and steep powerlaw (SPL) states (Ratheesh et al., 2024; Kushwaha et al., 2023; Rawat et al., 2023b; Rodriguez Cavero et al., 2023).
@@
$\bullet$ Cyg X-3: It was discovered over five decades ago Giacconi et al. (1967) and is a high-mass X-ray binary system hosting a compact object accreting from a Wolf-Rayet donor star van Kerkwijk et al. (1992). Despite extensive studies over the years, the nature of the compact object remains uncertain. Recent polarization measurements suggest that it could be a concealed Galactic ultra-luminous X-ray source (ULX) surpassing the Eddington limit due to anisotropic emissions Veledina et al. (2024a). Recent measurements from VLBA observations have precisely constrained the distance of the source as ${9.67}_{-0.48}^{+0.53}$ kpc Reid & Miller-Jones (2023).
@@
$\bullet$ LMC X-1: This is the first discovered extra-galactic persistent BH-XRB located in the Large Magellanic Cloud (LMC) Mark et al. (1969) at a well-constrained distance of $\sim 48.1$ kpc Pietrzyński et al. (2013). Dynamical measurements have determined the black hole’s mass to be $10.9\pm 1.4$ $M_{\odot}$ with a moderate inclination of $36.4^{{}^{\circ}}\pm 1.9^{{}^{\circ}}$ Orosz et al. (2009). The measurements using the continuum fitting method in the thermally dominated state constrain LMC X1-1 spin as $\sim 0.92$ Gou et al. (2009); Bhuvana et al. (2021), indicating the presence of a rapidly rotating black hole. LMC X-1 is the first source for which a null-detection of X-ray polarization has been reported from IXPE data (Podgorný et al., 2023).
$\bullet$ 4U 1957++115: This is a bright and persistent low-mass X-ray binary, first discovered by the Uhuru mission in $1973$ Giacconi et al. (1974). The source distance is found to be in the range of 5405-40 kpc, with inclination within $\sim 13^{\circ}-78^{\circ}$ and mass in $\sim 6-10M_{\odot}$ (Maitra et al., 2014; Gomez et al., 2015). Nevertheless, several studies suggest the presence of a maximally spinning black hole in 4U 1957++115 Barillier et al. (2023). Polarimetric study of the source reveals several relativistic effects close to the black hole, leading to low PD of $\sim 1.9\%$ (Marra et al., 2024).
@@
$\bullet$ LMC X-3: This is a persistent extra-galactic BH-XRB in Large Magellanic Cloud (LMC) at a distance of $48.1$ kpc Orosz et al. (2009), discovered by UHURU in $1971$. A slowly rotating black hole of spin $0.25_{-0.29}^{+0.20}$ Steiner et al. (2010); Bhuvana et al. (2021) and mass $6.98\pm 0.56~{\rm M}_{\odot}$ Orosz et al. (2014); Bhuvana et al. (2021) is found to be present at the central core of the system. LMC X3-3 is a relatively high inclined system ($i\sim 69.24^{\circ}\pm 0.72^{\circ}$) Orosz et al. (2014). Being a persistent source, it mostly remains in the HSS Bhuvana et al. (2021) except for a few occasions during which the LHS is also observed Smale & Boyd (2012). An exceptionally anomalous low state of peak luminosity $\sim 10^{35}$ erg $\rm s^{-1}$ is also observed for a few instances Torpin et al. (2017). Interestingly, LMC X-3 is another BH-XRB for which a relatively high polarization degree ($\sim 3\%$) has been reported from soft state observations (Majumder et al., 2024a; Svoboda et al., 2024a).